Maria is twice the age of jose right now. 30 years from now she will be 10 years older than jose and the sum of their ages will be 90 . what is maria and jose's ages now?

Answer by T-Maree(17)

Maria = 20 now
Jose = 10 now
20 + 30 = 50
10 + 30 = 40
50 + 40 = 90